The table below shows which hand is favored by each person from a sample of 100 people (50 male and 50 female). Use the information in the table to answer the following questions. Left-Handed Right-Handed Male 15 35 Female 10 40 Determine the probability that a randomly selected person from the sample will be female or left-handed. Number | % (Round to the nearest whole percentage) Compute a 95% confidence interval for the true proportion of the population that is left-handed. Number % to | Number | % Compute a 95% confidence interval for the true proportion of the population that is right-handed. Number % to | Number | %
Answer #1

A random sample of 100 people (50 males and 50 females) is selected and they are being asked if they use their left hand moreThe above function yields a critical value of 1.96. Substitute all the values in the above formula to compute the confidence0.75x(1-0.75) CI = 0.75-1.96, V 100 10.75x(1-0.75) = 0.75-1.96 100 10.75 0.25 = 0.75+1.96, 100 = 0.75+(1.96x0.0433) = 0.750.0